Calculate the volume of 0.250M solution made by dissolving 5.56g of solid magnesium chloride in water
1
Expert's answer
2019-05-02T07:28:53-0400

n(MgCl2) = m(MgCl2) / M(MgCl2) = 5.56 g / 95 g mol-1 = 0.0585 mol.

Vs(MgCl2) = n(MgCl2) / C(MgCl2) = 0.0585 mol / 0.250 M = 0.234 L = 234 mL.